# Quantize-Sample-and-Verify: LLM Acceleration via Adaptive Edge-Cloud Speculative Decoding

### v1

@@ Abstract
Index Terms:
Edge-cloud speculative decoding, large language models, reinforcement learning
@@ I Introduction
As illustrated in Fig. 1, edge-cloud speculative decoding (SD) employs a small language model (SLM) at the edge to generate multiple draft tokens, which are subsequently verified in parallel by a large language model (LLM) in the cloud [1, 2]. This setup exploits the low-latency benefits of parallel verification over traditional autoregressive generation, significantly improving token generation throughput[3, 4, 5, 6].
In edge-cloud deployments[7, 8], however, SD necessitates frequent communication between the SLM and the LLM, including the transmission of draft token sequences and their corresponding probability distributions. When operating over constrained or time-varying uplink channels, these communication demands can introduce nontrivial delays that hinder overall decoding performance.
@@
Edge–cloud collaborative SD was initially explored in [3], which proposed to leverage a lightweight on-device SLM to draft tokens and identify uncertain tokens for cloud-based verification by an LLM. To further reduce uplink transmissions to the cloud, the authors of the work [4] introduced an uncertainty-based skipping strategy, where tokens with high confidence are excluded from verification. In [5], the authors provided a case study about using SD on an image captioning task served at the edge. In parallel, the early exit technique was integrated into edge-cloud SD systems, enabling the client to preemptively draft subsequent tokens before final verification. This approach effectively utilizes idle time and enhances parallelism between the edge and cloud [6].
In prior work [3, 4] the quantization of the token distributions is performed after the sampling step following a sample-quantize (S-Q) approach. As a result, the distribution of the tokens generated via edge-cloud SD deviates from the token distribution that would have been produced via autoregressive generation from the cloud-based LLM. While in [5, 6], the token distributions need to be transmitted losslessly, which introduces heavy communication overload.
To address this problem, this letter proposes a novel quantize-sample (Q-S) strategy that strictly aligns the output distribution of edge-cloud SD with that of the LLM. Furthermore, rather than assuming negligible communication latency as in prior art, this letter explicitly models both uplink and downlink transmission delays, capturing the effects of dynamic transmission conditions. Based on this analysis, we introduce an adaptive resource allocation mechanism that dynamically tunes draft length and quantization precision based on both semantic uncertainty and current channel conditions.
@@ II-2 Uplink Transmission
Each token requires $\lceil{\log_{2}(V)}\rceil$ bits to represented losslessly, while the number of bits used to quantize each probability vector $q^{t}_{l}$ for $l=1,2,...,L^{t}$ is denoted as $b^{t}$. We denote the quantized probability vectors as $\mathbf{\hat{Q}}^{t}=[\mathbf{\hat{q}}_{1}^{t},\mathbf{\hat{q}}_{2}^{t},...,\mathbf{\hat{q}}_{L^{t}}^{t}]$, where the $l$-th probability vector is expressed as $\mathbf{\hat{q}}_{l}^{t}=[{\hat{q}}_{l,1}^{t},{\hat{q}}_{l,2}^{t},...,{\hat{q}}_{l,V}^{t}]^{T}$.
@@ II-3 Token Verification
After receiving the token sequence $\mathbf{x}^{t}$ and probabilities $\mathbf{\hat{Q}}^{t}$, the server employs an LLM for token verification. Note that, by design, the server is aware of all the previously generated tokens, and thus it has access to the prefix $\mathbf{s}^{t}$. Using the prefix and the received token sequence $\mathbf{x}^{t}$ as inputs, the server can produce the $L^{t}+1$ probability vectors $\mathbf{p}_{1}^{t},\mathbf{p}_{2}^{t},\dots,\mathbf{p}_{L^{t}}^{t},\mathbf{p}_{L^{t}+1}^{t}$ in parallel, where each probability vector $\mathbf{p}_{l}^{t}=[{p}_{l,1}^{t},{p}_{l,2}^{t},...,{p}_{l,V}^{t}]^{T}$ depends on the prefix $\mathbf{s}^{t}$ and on the tokens $[x_{1}^{t},x_{2}^{t},...,x_{l-1}^{t}]^{T}$. The server then uses the probability vector $\mathbf{p}_{l}^{t}$ to decide whether to accept or reject the $l$-th token $x_{l}^{t}$.
@@ II-4 Downlink Transmission
After the verification phase, the device needs to be notified about $N^{t}$, the number of tokens accepted by the server, as well as about the newly generated token at the server, namely $\tilde{x}^{t}_{l}$ or $\tilde{x}^{t}_{L^{t}+1}$. This requires the transmission of $\lceil\log_{2}(L^{t})\rceil+\lceil{\log_{2}(V)}\rceil$ bits.
@@ III-A Quantize-Sample-and-Verify
This result follows directly from the known properties of SD [1] by viewing the quantized distribution $\hat{\mathbf{Q}}^{t}$ as the generating probability distribution for the SLM. In fact, the SD protocol accepts and samples tokens, via (1) and (2), using the quantized distributions.
@@ III-B Lattice-Based Quantization
While the validity property in Proposition 1 holds for any quantization scheme at any resolution, in this paper we adopt the quantization algorithm proposed in [9]. This method compresses probability vectors more effectively than uniform quantization schemes [4] by exploiting the sparse structure of language model outputs. The algorithm restricts probability vectors to a lattice $Q_{\ell}$ that forms a discrete subset of the $V$-dimensional probability simplex. Specifically, we choose a positive integer $\ell$ such that each probability $q_{i}$ in the vector can be expressed as a rational number with denominator $\ell$, i.e., $q_{i}=o_{i}/\ell$, where $o_{i}$ are non-negative integers summing to $\ell$. The set of all such quantized probability vectors of dimension $V$ is defined as
@@ IV-A Latency Analysis
We denote by $T_{\text{SLM}}$ and $T_{\text{LLM}}$ the time required for the SLM and LLM, respectively, to generate a single token. Typically, one has the strict inequality $T_{\text{SLM}}\ll T_{\text{LLM}}$ since the LLM has far more parameters, leading to higher memory access and computation latency. For example, on an NVIDIA A100 GPU, a large $13$B-parameter model might take around $32$ ms per token, whereas a small $125$M-parameter model needs only about $5$ ms.
In addition to processing delays, we must account for communication delays between edge and cloud at each iteration $t$. Let $L^{t}$ be the number of draft tokens generated by the SLM and $b^{t}$ be the number of quantization bits used for each token’s distribution at iteration $t$. The uplink transmission time for sending these $L^{t}$ token probabilities and associated token identities can be modeled as
@@
where $V$ is the vocabulary size; $\lceil\log_{2}(V)\rceil$ accounts for the bits needed to identify each token; $b^{t}$ is the number of bits for the quantized probability; and $C^{t}_{u}$ is the uplink data rate (in bits/s) at iteration $t$. Similarly, the downlink time to return the verification results for $L^{t}$ tokens can be expressed as
@@
where $L^{t}T_{\text{SLM}}$ is the time for the SLM to generate $L^{t}$ draft tokens (since it generates tokens one by one); $T^{t}_{u}$ is the uplink delay to send those drafts; $T_{\text{LLM}}$ is the cloud verification time (since the LLM verifies $L^{t}$ tokens in parallel); and $T^{t}_{d}$ is the downlink time to receive the results. If the SD process runs for a total of $I$ iterations, the overall decoding latency is the sum over all iterations
@@ IV-B Problem Definition
The input state $s^{t}$ to the policy $\bm{\pi}$ at the beginning of iteration $t$ includes: (i) semantic information in the form of the prefix confidence vector $\mathbf{f}^{t}=[f_{1},f_{2},\ldots,f_{\Delta{t}}]$, where each $f_{i}$ is the probability assigned by the SLM to token $i$ in the current prefix $\mathbf{s}^{t}$; and (ii) connectivity information in the form of the current uplink channel rate $C^{t}_{u}$.
@@ IV-C Reinforcement Learning-Based Throughput Optimization
To address the problem (12), we adopt a reinforcement learning approach. To this end, we design a reward signal to provide the learning agent with immediate feedback on the quality of its action $a^{t}$. Ideally, we want the reward to reflect the token throughput achieved. A natural definition of the reward at the terminal iteration $\tau^{*}$ would the overall throughput. However, assigning a reward only at the end of the sequence would make learning difficult due to sparse feedback. Instead, we give a non-zero reward at each intermediate iteration that quantifies the contribution of that iteration to the final throughput.
Specifically, we define the immediate reward at iteration $t$ as the throughput gained in that iteration, i.e.,
@@
where $\mathbf{x}^{t}=[x_{1}^{t},x_{2}^{t},...,x_{L^{t}}^{t}]^{T}$ denotes the set of $L^{t}$ draft tokens sampled by the SLM at iteration $t$, $\hat{q}(\mathbf{x}^{t}|\mathbf{s}^{t})$ is the joint distribution of those tokens under the SLM’s (quantized) output, and $g(\mathbf{x}^{t})$ is the expected number of tokens in $\mathbf{x}^{t}$ that are accepted by the LLM. Given that the LLM’s verification process can be modeled as a sequence of Bernoulli trials, the function $g(\mathbf{x}^{t})$ can be expressed as
$g\left(\mathbf{x}^{t}\right)=\sum_{l=1}^{L^{t}}l\left(\prod_{j=1}^{l-1}\alpha_{j}^{t}\right)\left(1-\alpha_{l}^{t}\right)+\left(L^{t}+1\right)\prod_{j=1}^{L^{t}}\alpha_{j}^{t}.$ (15)
Using this reward signal, we adopt the double deep Q-Network (DDQN) method [11]. The Q-network is composed of multiple lightweight linear layers, and its training follows the standard procedures described in [11].
@@ V-A Setting
For numerical performance evaluation, we adopt a simulation environment modeling a time-varying uplink channel through a discrete-time two-state Markov model. In this model, the uplink alternates between a low-rate state with capacity $C_{\mathrm{u}}^{t}=40$ kbits/s and a high-rate state with capacity $C_{\mathrm{u}}^{t}=400$ kbits/s. These rates are typical of bandwidth-constrained low-rate systems such as LoRA and NB-IoT[12]. The state transitions are governed by transition probabilities between low- and high-capacity states, and vice versa, given by $p_{\textrm{low}\rightarrow\textrm{high}}$ and $p_{\textrm{high}\rightarrow\textrm{low}}$, respectively. We assume delay-free downlink transmission. We indirectly control the number of quantization bits per token, $b^{t}$ by adjusting the integer $\ell^{t}$ in (5). A larger integer $\ell^{t}$ yields a finer probability resolution, and correspondingly a larger number of bits $b^{t}$.
We use the OPT-13B language model as the cloud LLM and the smaller OPT-125M model as the edge SLM [13]. The task is abstractive text summarization on the CNN/DailyMail dataset[14]. All experiments are carried out on a Nvidia A100 40GB GPU for consistency, the runtime of the small model was adjusted by applying a scaling factor derived from benchmark comparisons with representative edge GPUs, RTX3060.
@@
Figure 4: Token throughput (tokens per second) achieved by different strategies as a function of the LLM sampling temperature. For both Q-S and S-Q, we set $(L^{t}=3,b^{t}=704)$.
@@ V-B Results
Fig. 3(a) shows the ROUGE-2 score, which measures the overlap of bigrams (two-word sequences) between generated text and reference text from the data set [15], as a function of the sampling temperatures of both SLM and LLM. The figure confirms that existing S-Q strategies fail to match the accuracy of the LLM, particularly at high temperatures. Fig. 3(b) verifies that the token entropy of all methods increases with temperature, confirming that higher temperatures introduce greater randomness and output variability, which is beneficial for scenarios that require high text diversity[14]. In contrast, the proposed Q-S approach can maintain the same token quality as the LLM across all sampling temperatures, with or without dynamic optimization.
Fig. 4 plots the achieved throughput (tokens per second) versus the sampling temperature. The figure demonstrates the advantages of SD over LLM in terms of token throughput. As shown in Fig. 3(a), this comes without loss of accuracy when deploying Q-S, while the use of SD via S-Q would entail a reduction in accuracy. When using Q-S, the figure confirms that the proposed dynamic optimization can significantly enhance the throughput across all values of the temperature. We observe that the proposed dynamic strategy consistently outperforms all static baselines, as well as the heuristic adaptive strategies described earlier across the entire temperature range. This demonstrates the efficacy of dynamically tuning the draft length and quantization precision in response to different semantic and channel conditions using a reinforcement learning-based approach.

### v2

@@ I Introduction
As illustrated in Fig. 1, edge-cloud speculative decoding (SD) employs a small language model (SLM) at the edge to generate multiple draft tokens, which are subsequently verified in parallel by a large language model (LLM) in the cloud [1]. This setup exploits the low-latency benefits of parallel verification over traditional autoregressive generation, significantly improving token generation throughput[2, 3, 4, 5]. Compared with cloud-only SD, edge-cloud SD enhances resilience in the presence of intermittent network or cloud availability[6], while reducing operational costs by decreasing redundant API calls[7], and by contributing to improved energy efficiency and system scalability[8].
In edge-cloud deployments[9], however, SD necessitates frequent communication between the SLM and the LLM, including the transmission of draft token sequences and their corresponding probability distributions. When operating over constrained or time-varying uplink channels, these communication demands can introduce nontrivial delays that hinder overall decoding performance.
@@
Edge–cloud collaborative SD was initially explored in [4], which proposed to leverage a lightweight on-device SLM to draft tokens and identify uncertain tokens for cloud-based verification by an LLM. To further reduce uplink transmissions to the cloud, the authors of the work [5] introduced an uncertainty-based skipping strategy, where tokens with high confidence are excluded from verification. In [2], the authors provided a case study about using SD on an image captioning task served at the edge. In parallel, the early-exit mechanism was integrated into edge-cloud SD systems, enabling the client to preemptively draft subsequent tokens before final verification [3].
In prior works [2, 3], the token distributions were assumed to be transmitted losslessly from edge to cloud. To reduce the communication overhead, the papers [4] and [5] proposed to quantize the token distributions after the sampling step at the edge. For such sample-quantize (S-Q) schemes, the distribution of the tokens generated at the edge deviates from the token distribution used for verification at the cloud. As a result, S-Q SD violates the key property of SD of preserving the cloud LLM token distribution.
In this context, this work introduces a quantize-sample (Q-S) protocol for edge-cloud SD that strictly aligns the output distribution of edge-cloud SD with that of the LLM. Furthermore, rather than assuming negligible communication latency as in prior work, this letter explicitly models both uplink and downlink transmission delays, capturing the effects of dynamic transmission conditions. Based on this analysis, we introduce an adaptive resource allocation mechanism that dynamically tunes draft length and quantization precision based on both semantic uncertainty and current channel conditions.
@@ II-2 Uplink Transmission
Each token requires $\lceil{\log_{2}(V)}\rceil$ bits to be represented losslessly, while the number of bits used to quantize each probability vector $q^{t}_{l}$ for $l=1,2,...,L^{t}$ is denoted as $b^{t}$. We denote the quantized probability vectors as $\mathbf{\hat{Q}}^{t}=[\mathbf{\hat{q}}_{1}^{t},\mathbf{\hat{q}}_{2}^{t},...,\mathbf{\hat{q}}_{L^{t}}^{t}]$, where the $l$-th probability vector is expressed as $\mathbf{\hat{q}}_{l}^{t}=[{\hat{q}}_{l,1}^{t},{\hat{q}}_{l,2}^{t},...,{\hat{q}}_{l,V}^{t}]^{T}$.
@@ II-3 Token Verification
After receiving the token sequence $\mathbf{x}^{t}$ and probabilities $\mathbf{\hat{Q}}^{t}$, the server employs an LLM for token verification. Note that, by design, the server is aware of all the previously generated tokens, and thus it has access to the prefix $\mathbf{s}^{t}$. Using the prefix and the received token sequence $\mathbf{x}^{t}$ as inputs, the server can produce the $L^{t}+1$ probability vectors $\mathbf{p}_{1}^{t},\mathbf{p}_{2}^{t},\dots,\mathbf{p}_{L^{t}}^{t},\mathbf{p}_{L^{t}+1}^{t}$ in parallel, where each probability vector $\mathbf{p}_{l}^{t}=[{p}_{l,1}^{t},{p}_{l,2}^{t},...,{p}_{l,V}^{t}]^{T}$ depends on the prefix $\mathbf{s}^{t}$ and on the tokens $x_{1}^{t},x_{2}^{t},...,x_{l-1}^{t}]^{T}$. The server then uses the probability vector $\mathbf{p}_{l}^{t}$ to decide whether to accept or reject the $l$-th token $x_{l}^{t}$.
@@ II-4 Downlink Transmission
After the verification phase, the device needs to be notified about $N^{t}$, the number of tokens accepted by the server, as well as about the newly generated token at the server, namely $\tilde{x}^{t}_{l}$ or $\tilde{x}^{t}_{L^{t}+1}$. This requires the transmission of $\lceil\log_{2}(L^{t})\rceil+\lceil{\log_{2}(V)}\rceil$ bits, where $\lceil\log_{2}(L^{t})\rceil$ is the number of bits used to describe the integer $N_{t}$, and $\lceil{\log_{2}(V)}\rceil$ is the number of bits required to represent the new token produced by the cloud.
@@ III-A Quantize-Sample-and-Verify
This result follows directly from the known properties of SD [7] by viewing the quantized distribution $\hat{\mathbf{Q}}^{t}$ as the generating probability distribution for the SLM. In fact, the SD protocol accepts and samples tokens, via (1) and (2), using the quantized distributions. This proposition establishes the equivalence between the outputs of the cloud LLM and of Q-S edge-cloud SD protocol. This result thus demonstrates that employing Q-S SD is, in terms of token quality, equivalent to using the cloud LLM. Consequently, unlike S-Q methods [4, 5], the proposed scheme does not introduce any quality degradation. The key reason for this critical difference between S-Q and Q-S is that the former uses two distinct distributions for sampling and verification, while the latter uses the quantized edge model’s distribution for both steps.
@@ III-B Lattice-Based Quantization
In general, the validity property stated in Proposition 1 holds for any quantization scheme at any resolution. In this paper, we adopt a state-of-the-art quantization algorithm for probability vectors, namely lattice-based quantization [10]. Note that the literature also reports quantization schemes that operate at the level of prompts [11, 12] or of the key-value cache [13]. However, these methods are not applicable to the problem of quantizing probability distributions.
Lattice-based quantization restricts probability vectors to a lattice $Q_{\ell}$ that forms a discrete subset of the $V$-dimensional probability simplex. Specifically, we choose a positive integer $\ell$ such that each probability $q_{i}$ in the vector can be expressed as a rational number with denominator $\ell$, i.e., $q_{i}=o_{i}/\ell$, where $o_{i}$ are non-negative integers summing to $\ell$. The set of all such quantized probability vectors of dimension $V$ is defined as
@@
Lattice-based quantization algorithm has a complexity of $\mathcal{O}(V\log(V))$. In practice, this adds only a delay of a few microseconds per batch on modern GPUs.
@@ IV-A Latency Analysis
We denote by $T_{\text{SLM}}$ and $T_{\text{LLM}}$ the time required for the SLM and LLM, respectively, to generate a single token. Typically, one has the strict inequality $T_{\text{SLM}}\ll T_{\text{LLM}}$ since the LLM has far more parameters, leading to higher memory access and computation latency. For example, on an NVIDIA A100 GPU, a large 1313B-parameter model might take around $32$ ms per token, whereas a small 125125M-parameter model needs only about $5$ ms.
In addition to processing delays, we must account for communication delays between edge and cloud at each iteration $t$. Let $L^{t}$ be the number of draft tokens generated by the SLM and $b^{t}$ be the number of quantization bits used for each token’s distribution at iteration $t$. We consider standard packet-level transmission, which reduces protocol overhead as compared to per-token transmission. The uplink transmission time for sending these $L^{t}$ token probabilities and associated token identities can be modeled as
@@
where $V$ is the vocabulary size; $\lceil\log_{2}(V)\rceil$ accounts for the bits needed to identify each token; $b^{t}$ is the number of bits for the quantized probability; and $C^{t}_{\mathrm{u}}$ is the uplink data rate (in bits/s) at iteration $t$. Similarly, the downlink time to return the verification results for $L^{t}$ tokens can be expressed as
@@
where $L^{t}T_{\text{SLM}}$ is the time for the SLM to generate $L^{t}$ draft tokens (since it generates tokens one by one); $T^{t}_{\mathrm{u}}$ is the uplink delay to send those drafts; $T_{\text{LLM}}$ is the cloud verification time (since the LLM verifies $L^{t}$ tokens in parallel); and $T^{t}_{\mathrm{d}}$ is the downlink time to receive the results. If the SD process runs for a total of $I$ iterations, the overall decoding latency is the sum over all iterations
@@ IV-B Problem Definition
The input state $s^{t}$ to the policy $\bm{\pi}$ at the beginning of iteration $t$ includes: (i) semantic information in the form of the prefix confidence vector $\mathbf{f}^{t}=[f_{1},f_{2},\ldots,f_{\Delta{t}}]$, where each $f_{i}$ is the probability assigned by the SLM to token $i$ in the current prefix $\mathbf{s}^{t}$, and of the mean confidence feature $\bar{f}^{t}=\frac{1}{\Delta{t}}\sum_{i=1}^{\Delta{t}}f_{i}$; and (ii) connectivity information in the form of the current uplink channel rate $C^{t}_{\mathrm{u}}$. By incorporating the uplink rate $C_{\mathrm{u}}^{t}$ into the decision-making process, the agent can adaptively adjust its action policy based on connectivity information.
@@ IV-C Reinforcement Learning-Based Throughput Optimization
To address the problem (12), we adopt a reinforcement learning approach. In particular, we design a reward signal to provide the learning agent with immediate feedback on the quality of its action $a^{t}$. Ideally, we want the reward to reflect the token throughput achieved. A natural definition of the reward at the terminal iteration $\tau^{*}$ would be the overall throughput. However, assigning a reward only at the end of the sequence would make learning difficult due to sparse feedback. Instead, we give a non-zero reward at each intermediate iteration that quantifies the contribution of that iteration to the final throughput. Specifically, we define the immediate reward at iteration $t$ as the throughput gained in that iteration, i.e.,
@@
where $\mathbf{x}^{t}=[x_{1}^{t},x_{2}^{t},...,x_{L^{t}}^{t}]^{T}$ denotes the set of $L^{t}$ draft tokens sampled by the SLM at iteration $t$, $\hat{q}(\mathbf{x}^{t}|\mathbf{s}^{t})$ is the joint distribution of those tokens under the SLM’s (quantized) output, and $g(\mathbf{x}^{t})$ is the expected number of tokens in $\mathbf{x}^{t}$ that are accepted by the LLM. Given that the LLM’s verification process can be modeled as a sequence of Bernoulli trials, the function $g(\mathbf{x}^{t})$ can be expressed as $g\left(\mathbf{x}^{t}\right)=\sum_{l=1}^{L^{t}}l\left(\prod_{j=1}^{l-1}\alpha_{j}^{t}\right)\left(1-\alpha_{l}^{t}\right)+\left(L^{t}+1\right)\prod_{j=1}^{L^{t}}\alpha_{j}^{t}.$
We adopt a double deep Q-Network (DDQN) model to implement the policy $\bm{\pi}$ [15]. The Q-network consists of several lightweight linear layers, and is trained offline following the standard procedure in [15].
@@ V-A Setting
We adopt a simulation environment modeling a time-varying uplink channel through a discrete-time two-state Markov model. In this model, the uplink alternates between a low-rate state and a high-rate state. The state transitions are governed by transition probabilities between low- and high-capacity states, and vice versa, given by $p_{\textrm{low}\rightarrow\textrm{high}}$ and $p_{\textrm{high}\rightarrow\textrm{low}}$, respectively. We assume delay-free downlink transmission. We indirectly control the number of quantization bits per token, $b^{t}$ by adjusting the integer $\ell^{t}$ in (5). A larger integer $\ell^{t}$ yields a finer probability resolution, and correspondingly a larger number of bits $b^{t}$.
We use the OPT-13B language model as the cloud LLM and the smaller OPT-125M model as the edge SLM [16], targeting the task of abstractive text summarization on the CNN/DailyMail dataset [17]. All experiments are conducted on an NVIDIA A100 40GB GPU for consistency. Inference is performed with a batch size of $1$, which aligns with standard practice in existing literature[4, 5, 2, 3]. Rather than directly deploying OPT-125M on an edge device, both models are executed on the A100 GPU to ensure controlled and reproducible measurements. Specifically, we introduce a scaling factor to approximate the runtime difference between the A100 and the representative edge GPUs. The factor is derived from their token generation speed ratio of the given hardware platforms under identical prompts, and the measured latency $T_{\mathrm{SLM}}$ is then divided by this factor to simulate realistic edge-side inference latency. Code is available at this link.
@@
Figure 4: Token throughput achieved by different strategies as a function of sampling temperature. (a) Low-rate regime; (b) High-rate regime.
@@ V-B Results
The left panel of Fig. 3 shows the ROUGE-2 score, which measures the overlap of bigrams (two-word sequences) between generated text and reference text from the data set, as a function of the sampling temperatures of both SLM and LLM. The figure confirms that existing S-Q strategies fail to match the accuracy of the LLM, particularly at high temperatures. The right panel verifies that the token entropy of all methods increases with temperature, confirming that higher temperatures introduce greater randomness and output variability, which is beneficial for scenarios that require high text diversity[17]. In contrast, the proposed Q-S approach can maintain the same token quality as the LLM across all sampling temperatures, with or without dynamic optimization.
Fig. 4 plots the achieved throughput (tokens per second) versus the sampling temperature. Specifically, Fig. 4(a) shows the results in a low-rate regime ($350$ kbps on average), which is typical in bandwidth-constrained low-rate systems such as NB-IoT[18] and satellite-oriented communications[19]. Fig. 4(b) shows the results in a higher rate regime ($4$ Mbps on average). The figure demonstrates the advantages of SD over LLM in terms of token throughput. As shown in Fig. 4(a), this comes without loss of accuracy when deploying Q-S, while the use of SD via S-Q entails a reduction in accuracy.
When using Q-S, Fig. 4 confirms that the proposed dynamic optimization can significantly enhance the throughput across all values of the temperature. Specifically, we observe that the proposed dynamic strategy consistently outperforms all static baselines, as well as the heuristic adaptive strategies described earlier across the entire temperature range. We also note that the absolute gains in high-rate regime is smaller than that in low-rate regime, since communication becomes less of a performance bottleneck.